Tina used 1/3 of her bus tokens, then lost 3/4 of the rest. This left her with 3 tokens. How many did she start with?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

(X-1/3X)1/4=3 OR (2/3)X(1/4)=3 OR (2/12)X=3 OR 2X=36 OR X=18
PROOF 18-(1/3)18=12 THEN 12-(3/4)12=12-9=3 TOKENS LEFT OUT OF THE 18 TINA STARTED WITH.

Answer by stanbon(48568)

Tina used 1/3 of her bus tokens, then lost 3/4 of the rest. This left her with 3 tokens. How many did she start with?
----------
Let "x" be the original number of tokens.
EQUATION:
x-(1/3)x - (3/4)(2/3)x=3
(2/3)x - (1/2)x=3
(1/6)x=3
x=18 (original number of tokens)
Cheers,
Stan H.